# GFM-RAG: Graph Foundation Model for Retrieval Augmented Generation

### v1

@@ Abstract
Retrieval-augmented generation (RAG) has proven effective in integrating knowledge into large language models (LLMs). However, conventional RAGs struggle to capture complex relationships between pieces of knowledge, limiting their performance in intricate reasoning that requires integrating knowledge from multiple sources. Recently, graph-enhanced retrieval augmented generation (GraphRAG) builds graph structure to explicitly model these relationships, enabling more effective and efficient retrievers. Nevertheless, its performance is still hindered by the noise and incompleteness within the graph structure. To address this, we introduce GFM-RAG, a novel graph foundation model (GFM) for retrieval augmented generation. GFM-RAG is powered by an innovative graph neural network that reasons over graph structure to capture complex query-knowledge relationships. The GFM with 8M parameters undergoes a two-stage training process on large-scale datasets, comprising 60 knowledge graphs with over 14M triples and 700k documents. This results in impressive performance and generalizability for GFM-RAG, making it the first graph foundation model applicable to unseen datasets for retrieval without any fine-tuning required. Extensive experiments on three multi-hop QA datasets and seven domain-specific RAG datasets demonstrate that GFM-RAG achieves state-of-the-art performance while maintaining efficiency and alignment with neural scaling laws, highlighting its potential for further improvement [^1] Code: https://github.com/RManLuo/gfm-rag.
Graph Foundation Model, Retrieval Augmented Generation, Graph-enhanced Retrieval Augmented Generation
[…]
@@ 1 Introduction
Figure 1: The overview framework of GFM-RAG. We feed the query and a constructed KG-index into the graph foundation model retriever to obtain relevant documents for LLM generation.
@@
Recent advancements in large language models (LLMs) (OpenAI, 2024a; Meta, 2024; Yang et al., 2024) have greatly propelled the evolution of natural language processing, positioning them as foundational models for artificial general intelligence (AGI). Despite the remarkable reasoning ability (OpenAI, 2024b), LLMs are still limited in accessing real-time information and lack of domain-specific knowledge, which is outside the pre-training corpus. To address these limitations, retrieval-augmented generation (RAG) (Gao et al., 2023) has become a popular paradigm in adding new knowledge to the static LLMs by retrieving relevant documents into the context of LLM generation.
@@
Existing RAG methods typically retrieve documents independently, making it difficult to capture complex relationships between pieces of knowledge (Karpukhin et al., 2020; Chen et al., 2023; Moreira et al., 2024). This limitation hampers the performance of LLMs in integrating knowledge across document boundaries, particularly in multi-hop reasoning tasks (Yang et al., 2018; Trivedi et al., 2022) and real-world applications like legal judgment (Kang et al., 2024) and medical diagnoses (Jin et al., 2019), which require reasoning over multiple sources. Although recent methods have expanded the retrieval process into multiple steps and incorporate LLM reasoning, they still encounter high computational costs due to iterative retrieval and reasoning with LLMs (Trivedi et al., 2023; Sun et al., 2024; Joshi et al., 2024).
@@
Recently, graph-enhanced retrieval augmented generation (GraphRAG) (Peng et al., 2024; Han et al., 2024) has emerged as a novel solution that builds a graph structure to explicitly model the intricate relationships between knowledge. This enables the development of a graph-enhanced retriever to identify relevant information using graphs. The structural nature of graphs allows GraphRAG to capture global context and dependencies among documents, significantly improving reasoning across multiple sources (Edge et al., 2024). Methods like HippoRAG (Gutiérrez et al., 2024) enhance retrieval by employing a personalized PageRank algorithm to locate relevant knowledge with graphs. However, these algorithms rely solely on the graph structure, which is often noisy or incomplete, limiting their overall performance. Alternative methods (Mavromatis & Karypis, 2024; He et al., 2024) incorporate graph neural networks (GNNs) into the retrieval process. These methods have shown impressive performance due to GNNs’ powerful reasoning capabilities in handling incomplete graphs (Galkin et al., 2024). Nevertheless, they still face limitations in generalizability since they require training from scratch on new datasets.
@@
Nowadays, the search for a foundation GNN model that can transfer and generalize across different datasets has been an active research topic. Ideally, a foundation GNN or graph foundation model (GFM) can benefit from large-scale training and generalize across diverse graphs (Mao et al., 2024; Liu et al., 2023). Efforts have been made to identify transferable graph tokens (e.g., motifs, sub-trees, and relation graphs) (Galkin et al., 2024; Wang et al., 2024; Xia et al., 2024) that can be shared among different graphs for GFM design. However, these methods primarily focus on graph-related tasks (e.g., node classification and link prediction), leaving the design of a GFM to enhance LLMs’ reasoning ability unexplored.
@@
To bridge the gap, in this paper, we propose an effective, efficient, and general graph foundation model for retrieval augmented generation (GFM-RAG), thereby enhancing LLMs’ reasoning ability. As shown in Figure 1, we create a knowledge graph index (KG-index) from documents in each dataset. The KG-index consists of interconnected factual triples pointing to the original documents, which serves as a structural knowledge index across multiple sources, enhancing the integration of diverse knowledge for complex reasoning tasks (Gutiérrez et al., 2024). Then, we present the graph foundation model retriever (GFM retriever), driven by a query-dependent GNN that captures complex query-knowledge relationships in a unified, transferable space of semantics and graph structure. Through multi-layer message passing, the GFM retriever enables efficient multi-hop retrieval in a single step, surpassing previous multi-step methods. The GFM retriever, with 8M parameters, undergoes a two-stage training: unsupervised KG completion pre-training and supervised document retrieval fine-tuning on large-scale datasets, including 60 knowledge graphs with over 14M triples and 700k documents. This large-scale training ensures the generalizability of GFM retriever to be applied directly to unseen datasets without further training.
@@
In experiments, GFM-RAG achieves state-of-the-art performance across three multi-hop QA datasets, demonstrating its effectiveness and efficiency in multi-hop reasoning. It also generalizes well across seven RAG datasets from diverse domains, such as biomedical, customer service, and general knowledge, without requiring additional training. Furthermore, GFM-RAG follows the neural scaling law (Hestness et al., 2017), whose performance benefits from training data and model size scaling, emphasizing its potential as a foundational model for future improvements.
The main contributions of this paper are as follows:
@@ 2 Related Work
Retrieval-augmented generation (RAG) (Gao et al., 2023) provides an effective way to integrate external knowledge into large language models (LLMs) by retrieving relevant documents to facilitate LLM generation. Early works adopt the pre-trained dense embedding model to encode documents as separate vectors (Karpukhin et al., 2020; Chen et al., 2023; Li et al., 2023b; Moreira et al., 2024), which are then retrieved by calculating the similarity to the query. Despite efficiency and generalizability, these methods struggle to capture complex document relationships. Subsequent studies have explored multi-step retrieval, where LLMs guide an iterative process to retrieve and reason over multiple documents (Trivedi et al., 2023; Jiang et al., 2023; Su et al., 2024). However, this approach is computationally expensive.
@@
Graph-enhanced retrieval augmented generation (GraphRAG) (Peng et al., 2024; Han et al., 2024) is a novel approach that builds graphs to explicitly model the complex relationships between knowledge, facilitating comprehensive retrieval and reasoning. Early research focuses on retrieving information from existing knowledge graphs (KGs), such as WikiData (Vrandečić & Krötzsch, 2014) and Freebase (Bollacker et al., 2008), by identifying relevant facts or reasoning paths (Li et al., 2023a; LUO et al., 2024). Recent studies have integrated documents with KGs to improve knowledge coverage and retrieval (Edge et al., 2024; Liang et al., 2024). A graph structure is built from these documents to aid in identifying relevant content for LLM generation (Dong et al., 2024). Based on graphs, LightRAG (Guo et al., 2024) incorporates graph structures into text indexing and retrieval, enabling efficient retrieval of entities and their relationships. HippoRAG (Gutiérrez et al., 2024) enhances multi-hop retrieval by using a personalized PageRank algorithm to locate relevant knowledge with graphs. However, the graph structure can be noisy and incomplete, leading to suboptimal performance. Efforts to incorporate GNNs into graph-enhanced RAG (Mavromatis & Karypis, 2024; He et al., 2024) have shown impressive results due to the strong graph reasoning capabilities of GNNs in handling incomplete graphs (Galkin et al., 2024). Nonetheless, these methods still limit in generalizability due to the lack of a graph foundational model.
@@
Graph Foundation models (GFM) aims to be a large-scale model that can generalize to various datasets (Mao et al., 2024; Liu et al., 2023). The main challenge in designing GFMs is identifying graph tokens that capture invariance across diverse graph data. For instance, ULTRA (Galkin et al., 2024) employs four fundamental relational interactions in knowledge graphs (KGs) to create a GFM for link prediction. OpenGraph (Xia et al., 2024) develops a graph tokenizer that converts graphs into a unified node token representation, enabling transformer-like GFMs for tasks such as link prediction and node classification. GFT (Wang et al., 2024) introduces a transferable tree vocabulary to construct a GFM that demonstrates effectiveness across various tasks and domains in graph learning. Despite these successful efforts, most methods primarily focus on conventional graph-related tasks. How to design a GFM to enhance the reasoning of LLM remains an open question.
@@ 3 Approach
In the first step, KG-index($\cdot$) constructs a knowledge graph index ${\mathcal{G}}$ from the document corpus ${\mathcal{D}}$, followed by our proposed graph foundation model retriever (GFM-Retriever), which is pre-trained on large-scale datasets. It retrieves top-$K$ documents based on any user query $q$ and knowledge graph index ${\mathcal{G}}$. The retrieved documents ${\mathcal{D}}^{K}$, along with the query $q$, are then input into a large language model (LLM) to generate the final answer $a$. These three main components in GFM-RAG are illustrated in Figure 2 and will be detailed next.
@@ 3.1 KG-index Construction
Conventional embedding-based index methods encode documents as separate vectors (Karpukhin et al., 2020; Chen et al., 2023; Moreira et al., 2024), which are limited in modeling the relationships between them. Knowledge graphs (KGs), on the other hand, explicitly capturing the relationships between millions of facts, can provide a structural index of knowledge across multiple documents (Edge et al., 2024; Gutiérrez et al., 2024). The structural nature of the KG-index aligns well with the human hippocampal memory indexing theory (Teyler & DiScenna, 1986), where the KG-index functions like an artificial hippocampus to store associations between knowledge memories, enhancing the integration of diverse knowledge for complex reasoning tasks (Gutiérrez et al., 2024).
To construct the KG-index, given a set of documents ${\mathcal{D}}$, we first extract entities ${\mathcal{E}}$ and relations ${\mathcal{R}}$ to form triples ${\mathcal{T}}$ from documents. Then, the entity to document inverted index $M\in\{0,1\}^{|{\mathcal{E}}|\times|{\mathcal{D}}|}$ is constructed to record the entities mentioned in each document. Such a process can be achieved by existing open information extraction (OpenIE) tools (Angeli et al., 2015; Zhou et al., 2022; Pai et al., 2024). To better capture the connection between knowledge, we further conduct the entity resolution (Gillick et al., 2019; Zeakis et al., 2023) to add additional edges ${\mathcal{T}}^{\texttt{+}}$ between entities with similar semantics, e.g., (USA, equivalent, United States of America). Therefore, the final KG-index ${\mathcal{G}}$ is constructed as ${\mathcal{G}}=\{(e,r,e^{\prime})\in{\mathcal{T}}\cup{\mathcal{T}}^{\texttt{+}}\}$. In implementation, we leverage an LLM (OpenAI, 2024a) as the OpenIE tool and a pre-trained dense embedding model (Santhanam et al., 2022) for entity resolution.
@@ 3.2 Graph Foundation Model (GFM) Retriever
The GFM retriever is designed to retrieve relevant documents based on any user query and the constructed KG-index. While the KG-index offers a structured representation of knowledge, it still suffers from incompleteness and noise, resulting in suboptimal retrieval performance when solely relying on its structure (Gutiérrez et al., 2024). Recently, graph neural networks (GNNs) (Wu et al., 2020) have shown impressive graph reasoning ability by capturing the complex relationships between knowledge for retrieval or question answering (Mavromatis & Karypis, 2024; He et al., 2024). However, existing GNNs are limited in generalizability, as they are usually trained on specific graphs (Mao et al., 2024; Liu et al., 2023), which limits their application to unseen corpora and KGs. Therefore, there is still a need for a graph foundation model that can be directly applied to unseen datasets and KGs without additional training.
@@ 3.2.1 Query-dependent GNN
Conventional GNNs (Gilmer et al., 2017) follow the message passing paradigm, which iteratively aggregates information from neighbors to update entity representations. Such a paradigm is not suitable for the GFM retriever as it is graph-specific and neglects the relevance of queries. Recent query-dependent GNNs (Zhu et al., 2021; Galkin et al., 2024) have shown promising results in capturing query-specific information and generalizability to unseen graphs, which is essential for the GFM retriever and can be formulated as:
@@
The query-dependent GNN exhibits better expressively (You et al., 2021) and logical reasoning ability (Qiu et al., 2024), which is selected as the backbone of our GFM retriever. It allows the GFM retriever to dynamically adjust the message passing process based on user queries and find the most relevant information on the graph.
@@
where $h^{l}_{e},h^{l}_{r}$ denote the entity and relation embeddings at layer $l$, respectively. The relation embeddings $h^{0}_{r}$ are also initialized using the same sentence embedding model as the query, reflecting their semantics (e.g., “born_in”), and updated by a layer-specific function $g^{l+1}(\cdot)$, implemented as a 2-layer MLP. The $\text{Msg}(\cdot)$ is operated on all triples in the KG to generate messages, which is implemented with a non-parametric DistMult (Yang et al., 2015) following the architecture of NBFNet (Zhu et al., 2021). For each entity, we aggregate the messages from its neighbors ${\mathcal{N}}_{r}(e)$ with relation $r$ using sum and update the entity representation with a single linear layer.

### v2

@@ Abstract
Retrieval-augmented generation (RAG) has proven effective in integrating knowledge into large language models (LLMs). However, conventional RAGs struggle to capture complex relationships between pieces of knowledge, limiting their performance in intricate reasoning that requires integrating knowledge from multiple sources. Recently, graph-enhanced retrieval augmented generation (GraphRAG) builds graph structure to explicitly model these relationships, enabling more effective and efficient retrievers. Nevertheless, its performance is still hindered by the noise and incompleteness within the graph structure. To address this, we introduce GFM-RAG, a novel graph foundation model (GFM) for retrieval augmented generation. GFM-RAG is powered by an innovative graph neural network that reasons over graph structure to capture complex query-knowledge relationships. The GFM with 8M parameters undergoes a two-stage training process on large-scale datasets, comprising 60 knowledge graphs with over 14M triples and 700k documents. This results in impressive performance and generalizability for GFM-RAG, making it the first graph foundation model applicable to unseen datasets for retrieval without any domain-specific fine-tuning required. Extensive experiments on three multi-hop QA datasets and seven domain-specific RAG datasets demonstrate that GFM-RAG achieves state-of-the-art performance while maintaining efficiency and alignment with neural scaling laws, highlighting its potential for further improvement.
@@ 1 Introduction
Recent advancements in large language models (LLMs) [47, 42, 70] have greatly propelled the evolution of natural language processing, positioning them as foundational models for artificial general intelligence (AGI). Despite the remarkable reasoning ability [48], LLMs are still limited in accessing real-time information and lack of domain-specific knowledge, which is outside the pre-training corpus. To address these limitations, retrieval-augmented generation (RAG) [12] has become a popular paradigm in adding new knowledge to the static LLMs by retrieving relevant documents into the context of LLM generation.
@@
Existing RAG methods typically retrieve documents independently, making it difficult to capture complex relationships between pieces of knowledge [30, 5, 43]. This limitation hampers the performance of LLMs in integrating knowledge across document boundaries, particularly in multi-hop reasoning tasks [72, 63] and real-world applications like legal judgment [28] and medical diagnoses [25], which require reasoning over multiple sources. Although recent methods have expanded the retrieval process into multiple steps and incorporate LLM reasoning, they still encounter high computational costs due to iterative retrieval and reasoning with LLMs [64, 59, 26].
@@
Recently, graph-enhanced retrieval augmented generation (GraphRAG) [51, 17] has emerged as a novel solution that builds a graph structure to explicitly model the intricate relationships between knowledge. This enables the development of a graph-enhanced retriever to identify relevant information using graphs. The structural nature of graphs allows GraphRAG to capture global context and dependencies among documents, significantly improving reasoning across multiple sources [9]. Methods like HippoRAG [16] enhance retrieval by employing a personalized PageRank algorithm to locate relevant knowledge with graphs. However, these algorithms rely solely on the graph structure, which is often noisy or incomplete, limiting their overall performance. Alternative methods [41, 18] incorporate graph neural networks (GNNs) into the retrieval process. These methods have shown impressive performance due to GNNs’ powerful multi-hop reasoning capabilities on graphs [73]. Nevertheless, they still face limitations in generalizability since they require training from scratch on new datasets.
@@
Nowadays, the search for a foundation GNN model that can transfer and generalize across different datasets has been an active research topic. Ideally, a foundation GNN or graph foundation model (GFM) can benefit from large-scale training and generalize across diverse graphs [40, 37]. Efforts have been made to identify transferable graph tokens (e.g., motifs, sub-trees, and relation graphs) [11, 66, 68] that can be shared among different graphs for GFM design. However, these methods primarily focus on graph-related tasks (e.g., node classification and link prediction), leaving the design of a GFM to enhance LLMs’ reasoning ability unexplored.
Figure 1: The overview framework of GFM-RAG.
@@
To bridge the gap, in this paper, we propose an effective, efficient, and general graph foundation model for retrieval augmented generation (GFM-RAG), thereby enhancing LLMs’ reasoning ability. As shown in Figure˜1, we create a knowledge graph index (KG-index) from documents in each dataset. The KG-index consists of interconnected factual triples pointing to the original documents, which serves as a structural knowledge index across multiple sources, enhancing the integration of diverse knowledge for complex reasoning tasks [16]. Then, we present the graph foundation model retriever (GFM retriever), driven by a query-dependent GNN that captures complex query-knowledge relationships in a unified, transferable space of semantics and graph structure. Through multi-layer message passing, the GFM retriever enables efficient multi-hop retrieval in a single step, surpassing previous multi-step methods. The GFM retriever, with 8M parameters, undergoes a two-stage training: self-supervised KG completion pre-training and supervised document retrieval fine-tuning on large-scale datasets, including 60 knowledge graphs with over 14M triples and 700k documents. This large-scale training ensures the generalizability of GFM retriever to be applied to unseen datasets without further training.
@@
In experiments, GFM-RAG achieves state-of-the-art performance across three multi-hop QA datasets, demonstrating its effectiveness and efficiency in multi-hop reasoning. It also generalizes well across seven RAG datasets from diverse domains, such as biomedical, customer service, and general knowledge, without requiring additional training. Furthermore, GFM-RAG follows the neural scaling law [19], whose performance benefits from training data and model size scaling, emphasizing its potential as a foundational model for future improvements. The main contributions of this paper are as follows:
@@ 2 Related Work
Retrieval-augmented generation (RAG) [12] provides an effective way to integrate external knowledge into large language models (LLMs) by retrieving relevant documents to facilitate LLM generation. Early works adopt the pre-trained dense embedding model to encode documents as separate vectors [30, 5, 34, 43], which are then retrieved by calculating the similarity to the query. Despite efficiency and generalizability, these methods struggle to capture complex document relationships. Subsequent studies have explored multi-step retrieval, where LLMs guide an iterative process to retrieve and reason over multiple documents [64, 24, 58]. However, this approach is computationally expensive.
@@
Graph-enhanced retrieval augmented generation (GraphRAG) [51, 17] is a novel approach that builds graphs to explicitly model the complex relationships between knowledge, facilitating comprehensive retrieval and reasoning. Early research focuses on retrieving information from existing knowledge graphs (KGs), such as WikiData [65] and Freebase [3], by identifying relevant facts or reasoning paths [33, 38, 50]. Recent studies have integrated documents with KGs to improve knowledge coverage and retrieval [9, 35]. A graph structure is built from these documents to aid in identifying relevant content for LLM generation [8]. Based on graphs, LightRAG [15] incorporates graph structures into text indexing and retrieval, enabling efficient retrieval of entities and their relationships. HippoRAG [16] enhances multi-hop retrieval by using a personalized PageRank algorithm to locate relevant knowledge with graphs. However, the graph structure can be noisy and incomplete, leading to suboptimal performance. Efforts to incorporate GNNs into graph-enhanced RAG [41, 18] have shown impressive results due to the multi-hop graph reasoning capabilities of GNNs in handling incomplete graphs [73]. Nonetheless, these methods still limit in generalizability due to the lack of a graph foundational model.
@@
Graph Foundation models (GFM) aims to be a large-scale model that can generalize to various datasets [40, 37]. The main challenge in designing GFMs is identifying graph tokens that capture invariance across diverse graph data. For instance, ULTRA [11] employs four fundamental relational interactions in knowledge graphs (KGs) to create a GFM with 0.2M parameters for link prediction. OpenGraph [68] develops a graph tokenizer that converts graphs into a unified node token representation, enabling transformer-like GFMs for tasks such as link prediction and node classification. GFT [66] introduces a transferable tree vocabulary to construct a GFM that demonstrates effectiveness across various tasks and domains in graph learning. Despite these successful efforts, most methods primarily focus on conventional graph-related tasks, and transformer-like GFMs [61, 60] struggle with large-scale graphs and capture logical association [52]. How to design a GNN-based GFM to enhance the reasoning of LLM remains an open question.
@@ 3 Approach
In the first step, KG-index($\cdot$) constructs a knowledge graph index $\mathcal{G}$} from the document corpus $\mathcal{D}$}, followed by our proposed graph foundation model retriever (GFM-Retriever), which is pre-trained on large-scale datasets. It retrieves top-$K$ documents based on any user query $q$ and knowledge graph index $\mathcal{G}$}. The retrieved documents ${\mathcal{D}}^{K}$, along with the query $q$, are then input into a large language model (LLM) to generate the final answer $a$. These three main components in GFM-RAG are illustrated in Figure˜2 and will be detailed next.
@@ 3.1 KG-index Construction
Conventional embedding-based index methods encode documents as separate vectors [30, 5, 43], which are limited in modeling the relationships between them. Knowledge graphs (KGs), on the other hand, explicitly capturing the relationships between millions of facts, can provide a structural index of knowledge across multiple documents [9, 16]. The structural nature of the KG-index aligns well with the human hippocampal memory indexing theory [62], where the KG-index functions like an artificial hippocampus to store associations between knowledge memories, enhancing the integration of diverse knowledge for complex reasoning tasks [16].
[…]
@@ 3.2 Graph Foundation Model (GFM) Retriever
The GFM retriever is designed to retrieve relevant documents based on any user query and the constructed KG-index. While the KG-index offers a structured representation of knowledge, it still suffers from incompleteness and noise, resulting in suboptimal retrieval performance when solely relying on its structure [16]. Recently, graph neural networks (GNNs) [67] have shown impressive multi-hop reasoning ability by capturing the complex relationships between knowledge for retrieval or question answering [41, 18]. However, existing GNNs are limited in generalizability, as they are usually trained on specific graphs [40, 37], which limits their application to unseen corpora and KGs. Therefore, there is still a need for a graph foundation model that can be directly applied to unseen datasets and KGs without additional training.
@@ 3.2.1 Query-dependent GNN
Conventional GNNs [14] follow the message passing paradigm, which iteratively aggregates information from neighbors to update entity representations. Such a paradigm is not suitable for the GFM retriever as it is graph-specific and neglects the relevance of queries. Recent query-dependent GNNs [78, 11] have shown promising results in capturing query-specific information and generalizability to unseen graphs, which is essential for the GFM retriever and can be formulated as:
@@
The query-dependent GNN is theoretically proven to exhibit multi-hop logical reasoning ability [21, 73, 52] (detailed in Appendix˜A), which is selected as the backbone of our GFM retriever. It allows the GFM retriever to dynamically adjust the message passing process based on user queries and find the most relevant information on the graph with multi-hop reasoning. The path interpretation for this multi-hop reasoning process is shown in Section˜4.8.
@@
$h_{e}^{l+1}=\text{Update}(h_{e}^{l},\text{Agg}(\{m_{e^{\prime}}^{l+1}|e^{\prime}\in{\mathcal{N}}_{r}(e),r\in{\mathcal{R}}\})),$ (9)
where $h^{l}_{e},h^{l}_{r}$ denote the entity and relation embeddings at layer $l$, respectively. The relation embeddings $h^{0}_{r}$ are also initialized using the same sentence embedding model as the query, reflecting their semantics (e.g., “born_in”), and updated by a layer-specific function $g^{l+1}(\cdot)$, implemented as a 2-layer MLP. The $\text{Msg}(\cdot)$ is operated on all triples in the KG to generate messages, which is implemented with a non-parametric DistMult [71] following the architecture of NBFNet [78]. For each entity, we aggregate the messages from its neighbors ${\mathcal{N}}_{r}(e)$ with relation $r$ using sum and update the entity representation with a single linear layer.